OpenStudy (anonymous):

Help with two multiple choice questions please? #QuickHelp [will MEDAL and FAN]

OpenStudy (anonymous):

@mathstudent55 quick help?

hartnn (hartnn):

Do you know that if x-1 is a factor (or equivalently x= 1 is a root), then sum of all the co-efficients will be = 0 ?

hartnn (hartnn):

similarly, if x+1 is the factor, then sum of co-efficients of odd powers of x equals sum of co-efficient of even powers of x

hartnn (hartnn):

example : x^2 -2x+1 x-1 is a factor, sum of all co-efficients = 1-2+1 = 0 x^2 +2x+1 x+1 is a factor, sum of co-efficients of even powers of x = 1+1 = 2 sum of co-efficients of odd powers of x = 2

hartnn (hartnn):

Apply these to your options of 1st question and you'll be able to eliminate 3 of them! :)

OpenStudy (anonymous):

Sorry I had left for a moment but thanks! Lets move onto the second one.

hartnn (hartnn):

sure, 2nd one is easier if \((x-n)^p\) is a factor, we say that 'n' is a root with the multiplicity of 'p'

hartnn (hartnn):

example: if we say (x+9)^5 is a factor, then -9 is a root with multiplicity of 5.

OpenStudy (anonymous):

I got D for the second question

hartnn (hartnn):

correct! :)

OpenStudy (anonymous):

Ok wait lol I'm still un sure about the first one

hartnn (hartnn):

what is the sum of all the co-efficients for the 1st option?

OpenStudy (anonymous):

6

hartnn (hartnn):

1-6 + 7+6 -8 that is the sum...how did you get 6?

OpenStudy (anonymous):

You meant first as in A? I see now.

hartnn (hartnn):

yeah, 1st option

OpenStudy (anonymous):

10

hartnn (hartnn):

1-6 + 7+6 -8 is actually = 0 quickly do the same for other choices.

hartnn (hartnn):

except for the 3rd. as there are 4 roots, the degree will be 4. but the 3rd optons has degree = 3. eliminated

OpenStudy (anonymous):

Would it be B?

hartnn (hartnn):

nopes, for b the sum isn't 0 also for d. so x=1 could only be the root of A.

OpenStudy (anonymous):

Oh ok I see now, thanks for your time!

hartnn (hartnn):

welcome ^_^